OpenStudy (anonymous):

Solve for the variable in this equation. 3(x+2/5)=x-3/1 x=-2/3 or x=-1and 7/30 or x=-5/6 or x=-23/30

OpenStudy (thomas5267):

Can you express the equation more clearly? What is x-3/1? Why divided by 1?

OpenStudy (anonymous):

I dont know its just what it says

OpenStudy (thomas5267):

Hmm... Interesting... \[ \begin{align*} 3\left(x+\frac{2}{5}\right)&=x-3 \\ 3x+\frac{6}{5}&=x-3 \\ 3x-x&=-3-\frac{6}{5} \\ 2x&=-\frac{15}{5}-\frac{6}{5} \\ 2x&=-\frac{21}{5} \\ x&=-\frac{21}{10} \end{align*} \] @PaigeMya

OpenStudy (anonymous):

Whats interesting?

OpenStudy (thomas5267):

Not a clue on what does x-3/1 means? Both \(\frac{x-3}{1} \text{ and } x-\frac{3}{1}\) means the same thing...